Suppose the investment demand and private saving supply curves in the market for loanable funds are
given by the following equations: I = 2000 – 100r S = 500 + 100r
where r represents the real interest rate in percentage points (eg. 10% is represented by 10), and
quantities are in billions. Assume that a closed economy and that initially the government is running a
balanced budget (ie. government saving initially equals 0).
(c) Return to the original supply and demand conditions. Suppose now that the government changes
policy and will run deficits equal to 500 to finance current government services. How does this
affect the level of national saving? Calculate the new equilibrium. What is the level of private
saving in the new equilibrium?
(d) With reference to an aggregate production function, explain why the deficit described above might
affect future economic growth.
1
Expert's answer
2015-03-12T10:31:12-0400
I = 2000 – 100r
S = 500 + 100r
(c) BS = -$500 billion.
In this case the level of national saving will decrease. The new equilibrium will be:
I = S = 1250 - 500 = $750 billion
r = 750/100 - 5 = %2.5
(d) The deficit described above might decrease future economic growth
level, because there will be less saving and investment in new
production, so the increase in output will be lower or even negative.